# Generalizable Self-supervised Monocular Depth Estimation with Mixture of Low-Rank Experts for Diverse Endoscopic Scenes

### v1

@@ 4.2 Datasets
AF-SfMLearner[8] 0.168 4.440 13.870 0.204 0.770 0.105 1.772 11.590 0.134 0.889
MonoViT [6] 0.193 10.512 18.028 0.220 0.769 0.103 1.566 11.482 0.136 0.895
LiteMono [7] 0.179 6.366 15.196 0.216 0.754 0.124 2.314 13.156 0.175 0.820
Depth Anything[29] 0.154 3.616 12.733 0.189 0.784 0.082 1.122 9.409 0.110 0.929
Depth Anything v2[30] 0.182 4.994 15.067 0.219 0.740 0.092 1.328 10.407 0.120 0.918
IID-SfMLearner[9] 0.171 4.526 14.066 0.206 0.767 0.112 1.956 12.193 0.138 0.878
DVSMono[27] 0.143 2.956 11.905 0.181 0.796 0.095 1.412 10.811 0.127 0.907
DARES[12] 0.164 4.456 13.625 0.201 0.783 0.084 1.179 9.932 0.113 0.930
EndoDAC[8] 0.156 3.854 12.936 0.193 0.791 0.084 1.223 9.683 0.113 0.922
MonoPCC[28] 0.158 3.889 13.205 0.194 0.782 0.091 1.265 10.123 0.117 0.915
GD-EMoDE 0.140 3.092 11.749 0.175 0.808 0.069 0.730 7.597 0.088 0.963
4.3 Depth Estimation Results
4.3.1 Results on Realistic Data
The proposed method is evaluated on three different realistic datasets, SCARED, Hamlyn and SERV-CT. Meanwhile, our method is compared with state-of-the-art works including methods for endoscopic scenes (MonoPCC[28], EndoDAC[11], DVSMono[27], DARES[12], SfM-Diffusion[35], IID-SfM[9], Surgical-DINO[10], AF-SfM[8] and method from [4]) and natural scenes (MonoViT[6], LiteMono[7] and Depth Anything[29, 30]). All frameworks are trained and tested on SCARED dataset at first, results of which are shown in Table. 1 and Fig. 6. To evaluate the generalizability of the model, all frameworks trained on SCARED dataset are utilized to perform zero-shot depth estimation on Hamlyn and SERV-CT dataset, and the experimental results are shown in Table. 2 and Fig. 7. Note that some works without published codes and weights, including Surgical-DINO[10], Yang’s method[4] and SfM-Diffusion[35], are not compared for zero-shot estimation.
@@ 4.3.2 Results on Simulated Data and Sim-to-Real
AF-SfMLearner[8] 0.086 0.080 0.426 0.109 0.950 0.142 1.495 7.897 0.188 0.804
IID-SfMLearner[9] 0.080 0.069 0.429 0.108 0.959 0.140 1.484 7.733 0.186 0.810
EndoDAC[11] 0.107 0.110 0.530 0.133 0.924 0.103 0.652 5.306 0.127 0.918
Ours 0.078 0.056 0.421 0.109 0.959 0.096 0.635 5.241 0.120 0.923
@@
4.4 Ablation Studies
4.4.1 Ablations on Parameter-efficient Finetuning
Ablation studies are performed to demonstrate the effect of the proposed parameter-efficient finetuning based on block-wise mixture of low-rank experts (BW-MoLE), as Table 4 shows. Firstly, the allocation of the experts are replaced with group-wise allocation (GW-MoLE) and average allocation. The number of experts are set to 2, 4, 6 and 8 for each group consisting of four Transformer blocks for group-wise allocation, while the number of experts in all blocks are equal to x, represented as MoLE(top-k/x), for average allocation. Note that the deeper group is also allocated with more experts in GW-MoLE due to Fig. 7. Furthermore, the finetuning of each linear layer is demonstrated based on ablation studies. At last, the finetuning based on the proposed method is also compared with the PEFT with multiple kinds of LoRA.
4.4.2 Ablations on Self-supervised Learning
Ablation studies are performed to demonstrate the effect of the proposed self-supervised training strategy, as Table 5 shows. For our training strategy, the experiments are implemented to prove the effect of the loss from augmented image, Retinex algorithm and illumination-free alignment. Moreover, the proposed model is also trained based on the existing training pipeline AF-SfMLearner [8] and IID-SfMLearner [36] for comparison.
@@ 4.4.2 Ablations on Self-supervised Learning
HydraLoRA[37] 0.052 / 0.098 / 0.155 0.361 / 1.675 / 4.066 4.525 / 10.931 / 12.986 0.073 / 0.125 / 0.191 0.982 / 0.904 / 0.796
AdaLoRA[38] 0.054 / 0.076 / 0.144 0.438 / 0.985 / 3.057 4.949 / 8.992 / 12.075 0.078 / 0.103 / 0.179 0.978 / 0.933 / 0.801
DV-LoRA[11] 0.050 / 0.086 / 0.151 0.334 / 1.172 / 3.639 4.342 / 9.403 / 12.469 0.071 / 0.108 / 0.186 0.983 / 0.929 / 0.798
LoRA[13] 0.049 / 0.081 / 0.150 0.336 / 1.065 / 3.681 4.381 / 9.035 / 12.508 0.070 / 0.103 / 0.185 0.983 / 0.945 / 0.800
@@
AF-SfMLearner[8] 0.049 / 0.083 / 0.142 0.339 / 1.071 / 3.138 4.390 / 9.086 / 11.842 0.070 / 0.107 / 0.180 0.983 / 0.932 / 0.805
IID-SfMLearner[9] 0.058 / 0.092 / 0.148 0.448 / 1.342 / 3.302 4.997 / 9.924 / 12.372 0.081 / 0.118 / 0.184 0.976 / 0.913 / 0.794
@@
4.5 3D Reconstruction Results
3D reconstruction is a significant application of depth estimation for endoscopy. Therefore, we also perform 3D scene reconstruction based on the estimated depth maps using Open3D (https://www.open3d.org/). As Fig. 9 shows, the reconstructed scenes based on the depth maps provided by our method is with more clear and accuracy texture and appearance, compared with results from [11] and [28].
@@
AF-SfMLearner[8] 0.0941 0.0841 0.0742 0.0682
Depth Anything[29] 0.0932 0.0815 0.0794 0.0696
IID-SfMLearner[9] 0.0951 0.0851 0.0764 0.0682
EndoDAC[11] 0.0936 0.0832 0.0776 0.0687
MonoPCC[28] 0.1040 0.0842 0.0781 0.0939
Ours 0.0927 0.0814 0.0754 0.0678
4.6 Ego-motion Estimation Results
Ego-motion estimation, as one of important part in the self-supervised framework, is also evaluated with experiments. Four sequences from the test split of SCARED dataset are selected to perform ego-motion estimation. The accuracy of ego-motion estimation is evaluated and compared based on Absolute Trajectory Error (ATE), following previous works. Based on experimental results in Table 6 and Fig. 10, the proposed method can perform the most accurate ego-motion estimation, compared with multiple existing methods including MonoPCC[28], EndoDAC[11], IID-SfM[9] and AF-SfM[8].
@@ 5.1 Outstanding Accuracy and Generalization
Figure 11: Visualization of the experts’ weights based on t-SNE (Top) and qualitative comparison of results with different SSL pipelines (Bottom).
@@ 5.3 Generalization on Diverse Illumination
Different from previous self-supervised training, both of the interference of reflectance and the brightness inconsistency are jointly dealt with by the proposed training framework. The superiority of our pipeline is proved based on qualitative results from ablation studies and comparison with existing pipelines. Furthermore, the proposed self-supervised training strategy is compared with previous pipelines on the sample from SCARED dataset in Fig. 11, which demonstrates the filtering of influence from light inconsistency by our method.

### v2

@@ 4.3.1 Results on Realistic Data
AF-SfMLearner(Shao et al., 2022) 0.168 4.440 13.870 0.204 0.770 0.105 1.772 11.590 0.134 0.889
MonoViT (Zhao et al., 2022) 0.193 10.512 18.028 0.220 0.769 0.103 1.566 11.482 0.136 0.895
LiteMono (Zhang et al., 2023a) 0.179 6.366 15.196 0.216 0.754 0.124 2.314 13.156 0.175 0.820
Depth Anything(Yang et al., 2024a) 0.154 3.616 12.733 0.189 0.784 0.082 1.122 9.409 0.110 0.929
Depth Anything v2(Yang et al., 2024b) 0.182 4.994 15.067 0.219 0.740 0.092 1.328 10.407 0.120 0.918
IID-SfMLearner(Li et al., 2024) 0.171 4.526 14.066 0.206 0.767 0.112 1.956 12.193 0.138 0.878
DVSMono(Zhou et al., 2024) 0.143 2.956 11.905 0.181 0.796 0.095 1.412 10.811 0.127 0.907
DARES(Zeinoddin et al., 2024) 0.164 4.456 13.625 0.201 0.783 0.084 1.179 9.932 0.113 0.930
EndoDAC(Shao et al., 2022) 0.156 3.854 12.936 0.193 0.791 0.084 1.223 9.683 0.113 0.922
MonoPCC(Wang et al., 2025) 0.158 3.889 13.205 0.194 0.782 0.091 1.265 10.123 0.117 0.915
[…]
Figure 7: Qualitative results of zero-shot depth estimation on Hamlyn and SERV-CT dataset. The error maps denotes the relative error distribution in the area marked by the green box. Note that if there is no green box, the error map denotes the errors in the whole picture. The image below the input raw image is the ground truth of depth map in corresponding area.
@@ 4.3.2 Results on Simulated Data and Sim-to-Real
AF-SfMLearner(Shao et al., 2022) 0.086 0.080 0.426 0.109 0.950 0.142 1.495 7.897 0.188 0.804
IID-SfMLearner(Li et al., 2024) 0.080 0.069 0.429 0.108 0.959 0.140 1.484 7.733 0.186 0.810
EndoDAC(Cui et al., 2024b) 0.107 0.110 0.530 0.133 0.924 0.103 0.652 5.306 0.127 0.918
[…]
@@
HydraLoRA(Tian et al., 2024) 0.052 / 0.098 / 0.155 0.361 / 1.675 / 4.066 4.525 / 10.931 / 12.986 0.073 / 0.125 / 0.191 0.982 / 0.904 / 0.796
AdaLoRA(Zhang et al., 2023b) 0.054 / 0.076 / 0.144 0.438 / 0.985 / 3.057 4.949 / 8.992 / 12.075 0.078 / 0.103 / 0.179 0.978 / 0.933 / 0.801
DV-LoRA(Cui et al., 2024b) 0.050 / 0.086 / 0.151 0.334 / 1.172 / 3.639 4.342 / 9.403 / 12.469 0.071 / 0.108 / 0.186 0.983 / 0.929 / 0.798
LoRA(Hu et al., 2022) 0.049 / 0.081 / 0.150 0.336 / 1.065 / 3.681 4.381 / 9.035 / 12.508 0.070 / 0.103 / 0.185 0.983 / 0.945 / 0.800
4.4 Ablation Studies
4.4.1 Ablations on Parameter-efficient Finetuning
Ablation studies are performed to demonstrate the effect of the proposed parameter-efficient finetuning based on block-wise mixture of low-rank experts (BW-MoLE), as Table 4.3.2 shows. Firstly, the allocation of the experts are replaced with group-wise allocation (GW-MoLE(Gao et al., 2025)) and average allocation. Note that the deeper group is also allocated with more experts in GW-MoLE according to Fig. 7. The numbers of experts are set to 2, 4, 6 and 8 for each group consisting of four Transformer blocks for group-wise allocation, while the top-ks in all blocks are equal to 2. Moreover, the number of experts is set to x with the top-k setting, represented as MoLE(k/x), for average allocation. Furthermore, the finetuning of each linear layer is demonstrated based on ablation studies. At last, the finetuning based on the proposed method is also compared with the PEFT with multiple kinds of LoRA.
4.4.2 Ablations on Self-supervised Learning
Ablation studies are performed to demonstrate the effect of the proposed self-supervised training strategy, as Table 5 shows. For our training strategy, the experiments are implemented to prove the effect of the loss from augmented image, Retinex algorithm and illumination-free alignment. Moreover, the proposed model is also trained based on the existing training pipeline AF-SfMLearner (Shao et al., 2022) and IID-SfMLearner (Liu et al., 2020) for comparison.
@@ 4.4.2 Ablations on Self-supervised Learning
AF-SfMLearner(Shao et al., 2022) 0.049 / 0.083 / 0.142 0.339 / 1.071 / 3.138 4.390 / 9.086 / 11.842 0.070 / 0.107 / 0.180 0.983 / 0.932 / 0.805
IID-SfMLearner(Li et al., 2024) 0.058 / 0.092 / 0.148 0.448 / 1.342 / 3.302 4.997 / 9.924 / 12.372 0.081 / 0.118 / 0.184 0.976 / 0.913 / 0.794
4.5 3D Reconstruction Results
3D reconstruction is a significant application of depth estimation for endoscopy. Therefore, we also perform 3D scene reconstruction based on the estimated depth maps using Open3D  [^3] https://www.open3d.org/. As Fig. 9 shows, the reconstructed scenes based on the depth maps provided by our method is with more clear and accuracy texture and appearance, compared with results from (Cui et al., 2024b) and (Wang et al., 2025).
@@
4.6 Ego-motion Estimation Results
Ego-motion estimation, as one of important part in the self-supervised framework, is also evaluated with experiments. Four sequences from the test split of SCARED dataset are selected to perform ego-motion estimation. The accuracy of ego-motion estimation is evaluated and compared based on Absolute Trajectory Error (ATE), following previous works. Based on experimental results in Table 6 and Fig. 10, the proposed method can perform the most accurate ego-motion estimation, compared with multiple existing methods including MonoPCC(Wang et al., 2025), EndoDAC(Cui et al., 2024b), IID-SfM(Li et al., 2024) and AF-SfM(Shao et al., 2022).
@@ 4.6 Ego-motion Estimation Results
AF-SfMLearner(Shao et al., 2022) 0.0941 0.0841 0.0742 0.0682
Depth Anything(Yang et al., 2024a) 0.0932 0.0815 0.0794 0.0696
IID-SfMLearner(Li et al., 2024) 0.0951 0.0851 0.0764 0.0682
EndoDAC(Cui et al., 2024b) 0.0936 0.0832 0.0776 0.0687
MonoPCC(Wang et al., 2025) 0.1040 0.0842 0.0781 0.0939
[…]
@@ 5.1 Outstanding Accuracy and Generalization
Figure 11: Visualization of the experts’ weights based on t-SNE.
@@ 5.2 Generalization on Diverse Scenes
Figure 12: Comparison of results from models trained via different pipelines. The layout of this figure follows Fig.7.
@@ 5.3 Generalization on Diverse Illumination
Different from previous self-supervised training, both of the interference of reflectance and the brightness inconsistency are jointly dealt with by the proposed training framework. The superiority of our pipeline is proved based on qualitative results from ablation studies and comparison with existing pipelines. Furthermore, the proposed self-supervised training strategy is compared with previous pipelines on the sample from SCARED (first two rows) and Hamlyn (the last row) dataset in Fig. 12, which demonstrates the filtering of influence from light inconsistency by our method.
5.4 Limitations & Future Works
Although the proposed method achieves state-of-the-art performance compared with existing methods on monocular depth estimation in diverse endoscopic scenes, The inference speed of our method is lower than previous works while the speed still achieves about 30fps which could meet the real-time requirement, shown in Table.7. Although the number of training parameters remains low in the proposed framework, the step-by-step training strategy and the usage of foundation model still need lots of computational resources. In the future, more efficient training pipeline and foundation model could be developed for endoscopic depth estimation. Meanwhile, multimodal visual features such as semantic feature could boost the performance of depth estimation in endoscopy.
Table 7: Comparison of Inference Speed
Methods Speed(ms/frame)
SCARED Hamlyn
Depth Anything(Yang et al., 2024a, b) 5.0 3.0
IID-SfMLearner(Li et al., 2024) 2.0 1.0
DVSMono(Zhou et al., 2024) 12.7 10.5
EndoDAC(Cui et al., 2024b) 5.7 4.0
MonoPCC(Wang et al., 2025) 12.6 10.5
[…]